# Privacy Preserving Mechanisms for Coordinating Airspace Usage in Advanced Air Mobility

### v1

@@ 4.2 Step 2: Computing Integral Allocation
13:         end if
14:    end while
15:end for
Remark 4.2.
Note that in Algorithms 1-2, the SP does not require any information about the private valuations of individual AAM vehicles. Furthermore, each AAM vehicle is unable to infer the demand or private valuations of other vehicles in the airspace, ensuring that privacy is preserved. The iterative update in (5) can be interpreted as a learning mechanism, where AAM vehicles adapt to the environment based on signals received from the SP, progressively refining their demand through this distributed learning process.
5 Drone Delivery in Toulouse: A Case Study
In this section, we validate our proposed mechanism using a dataset of simulated package deliveries by Airbus, as shown in Fig. 3. Particularly, we use a package delivery scenario where four suppliers in Toulouse, France, have warehouses on the periphery of the city and make deliveries in locations randomly distributed around the city. In the Airbus dataset, the demand is generated using a Poisson process. Each demand simulates delivery operations, where a UAV departs from a specific vertiport, navigates through different sectors, and returns to its original departure vertiport. The trajectory of every UAV is generated using a high-fidelity Airbus trajectory simulator which includes takeoff, cruise, and landing phases. While in the current simulation, the UAV returns to its original departure vertiport, our algorithm can accommodate a return to a different location. Trajectories last approximately 5 minutes staggered at different departing times to simulate a realistic operational system. The entire simulation provides 4,000 seconds of simulated flight data. For our analysis, we divide the city into 12 cruising regions and 4 vertiports regions, as shown in Fig. 3.
Each flight has a desired 4D trajectory (three spatial dimensions with time as the fourth dimension). We assign a uniformly random utility to each flight between $150$ and $250$. The utility for each flight is highest when it can depart and arrive at its preferred times, with the utility decreasing by a factor of $\delta=0.95$ for each time step delay in departure time. We consider the value agents get from the outside option of $1$ and from dropping out $40$. The preferred menu for every agent is to stay parked for a small valuation, take the highest utility path, or delay for up to four steps.
We implement our mechanism (Algorithms 1-2) in three receding horizon steps of the Toulouse example, scheduling auctions at 5-minute intervals. Each time unit in our simulation represents 15 seconds; therefore, 20 units of time between horizon steps correspond to 5 minutes between rounds. Before the start, the SP assigns each agent a per-auction budget of air credits to all participating vehicles, between 150 and 250, which they have access to for each round. Before the start of every round, the SP gathers the demand for AAM vehicles requesting access to airspace. The overall budget of any UAV is comprised of the new air credits they received and any remaining budget from a previous round.
Figure 4: Evolution of prices and congestion of various sectors with time for the first three auctions of the Toulouse example. Only goods with non-zero prices are included. Regular lines, dashed lines, and dot-dashed lines denote the results of the 1st, 2nd, and 3rd auctions, respectively. The figure only includes goods where the price is non-zero at some timestamp. a) Price of goods vs time shows some positive prices for sectors 1, 2, and 4. b) Cumulative sector congestion shows the total allocated airspace after each auction, including airspace that was allocated in a previous auction. c) The price of arrival and departure slots shows positive prices for some arrival slots.
In Fig. 4, we present the evolution of prices and congestion across various sectors overtime during the first three auctions of the Toulouse example. Only the sectors with non-zero prices at some point are included, and we observe that capacity constraints are consistently met in all sectors. In the top plot, three sectors show positive prices in at least one of the auctions, with these prices corresponding to times of maximum sector utilization, as illustrated in the second plot [^8] Recall our market clearing definition: congestion can reach a maximum in a sector even if prices are zero.. An interesting observation in the first plot is that prices for sector S004 become positive only during the third auction, despite some drones being allocated to it for future use as early as the first auction (between 0 and 20 time units). Sector S004, however, only reaches maximum capacity during the third auction, resulting in the observed positive prices. Additionally, the prices for sectors S001 and S002 before and around time step 40 are considerably higher than any agent’s budget, which might initially appear inconsistent with five agents being allocated to these sectors at that time, as shown in the second plot. However, these agents were actually allocated in the first auction, when prices were low, as indicated by the solid line in the first plot. The subsequent high prices are reasonable because the sectors are at maximum capacity and cannot accommodate new requests in auctions 2 and 3. The third plot in Fig. 4 shows non-zero prices for certain departure and arrival slots. Overall, our mechanism effectively maintains congestion below the respective thresholds for all sectors, vertiports, and take-off/landing slots—at 5, 10, and 1—despite requests that exceed these capacities.
(a) Fixed point error per agent
(b) MCE
(c) Error on each agent’s linear constraints
(d) Error between each agent’s x and y demand
Figure 5: Convergence of Error Terms
In Fig. 5, we present the evolution of various errors resulting from Algorithm 1 in the first auction round. First, in Fig. 5(a), we analyze the evolution of the fixed point error, $\|\lambda^{(n,k)}-\omega^{(k)}\|$, as both $n$ and $k$ increase within Algorithm 1. Here, the inner loop length is set to $N=20$. The fixed point error consistently decreases with additional inner loop rounds, suggesting a contraction property in the fixed point map (FP). Further research could explore sufficient conditions that ensure this contraction in relation to the underlying AAM market parameters. In Fig. 5(b), we present the evolution of the market clearing error (MCE, defined in (9)), which ultimately approaches zero. This indicates that resources with positive prices achieve a balance between demand and supply, while zero-priced resources satisfy capacity constraints. In Fig. 5(c), we show the evolution of the individual constraint error (ICE, defined in (10)), which also converges to zero, indicating satisfaction of (4e). Notably, we observe spikes at the start of each inner loop due to updates in the budget adjustment parameter, which alter the outcome of (5). Lastly, in Fig. 5(d), we present the evolution of the baseline demand error (BDE, defined in (11)), which also converges to zero, confirming that (4b) is met. This analysis highlights Algorithm 1’s effectiveness in reducing errors and satisfying key constraints within the auction framework.
6 Vertiport Reservation Mechanism in Northern California
Figure 6: Northern California Vertiport Map. This map, based on an image from Google Maps, highlights seven distinct vertiports with unique color codes, showing the example routes in red lines for clarity.
In this example, 20 air taxis request a departure, air route, and landing clearances among seven vertiport destinations during a 10-minute auction window. The requests by the air taxis, the final allocation of routes, and payments to the SP are presented in Table 1. Additional information from the network in Table 2 displays the maximum capacity in every segment of the desired routes, the utility of the air taxis for a given path, and their initial air credits. We set $\beta=50,p_{\textsf{o}}=10,v_{u,\textsf{o}}=1,v_{u,\varnothing}=1,N=2,\textsf{tol}=1\times 10^{-4}$. In these tables, air taxis marked with the same color indicate that they demanded the same trajectory slot simultaneously, resulting in a constraint violation. Consequently, only a subset of these air taxis were allocated their preferred route, while others were denied access.
Below, we highlight the main observations from our numerical study.
@@ 6 Vertiport Reservation Mechanism in Northern California
Table 1: Allocation of air taxis to the desired route and the corresponding payments
Agent Req. Route Req. (Dep, Arr) Allocated (Dep, Arr) Status Payment (\$)
AC001 (V007, V002) (16, 54) (16, 54) on-time 0.0
AC002 (V005, V004) (19, 47) (19, 47) on-time 5.73
AC011 (V006, V004) (19, 47) (20, 48) delayed 1.53
AC003 (V002, V001) (16, 21) (19, 24) delayed 3.71
AC004 (V002, V001) (16, 21) (16, 21) on-time 20.36
AC015 (V002, V001) (16, 21) (20, 25) delayed 0.86
AC013 (V002, V006) (17, 41) (17, 41) on-time 11.11
AC018 (V002, V007) (18, 56) (18, 56) on-time 8.46
AC005 (V003, V002) (11, 19) (11, 19) on-time 0.0
AC006 (V005, V007) (18, 68) (18, 68) on-time 0.0
AC007 (V003, V002) (15, 23) (15, 23) on-time 0.0
AC008 (V007, V001) (12, 54) (12, 54) on-time 0.0
AC009 (V001, V005) (13, 34) (14, 35) delayed 2.10
AC010 (V001, V005) (13, 34) (13, 34) on-time 5.75
AC012 (V005, V001) (16, 37) (16, 37) on-time 0.0
AC014 (V001, V002) (11, 24) (11, 24) on-time 0.0
AC016 (V007, V005) (17, 67) (17, 67) on-time 0.0
AC017 (V004, V006) (16, 44) (16, 44) on-time 0.0
AC019 (V004, V002) (16, 35) (16, 35) on-time 0.0
AC020 (V003, V006) (16, 38) (16, 38) on-time 0.11
Table 2: Utility and air-credits for different air taxis, along with the capacity constraints on their route
Aircraft Routes Max. Capacity (Dep, Route, Arr) Initial Air Credits Utility Status Rank
AC001 (V007, V002) (2,4,1) 125 118 on-time 6
AC002 (V005, V004) (4,5,1) 90 171 on-time 7
AC011 (V006, V004) (1,2,1) 78 135 delayed 19
AC003 (V002, V001) (1,1,2) 135 172 delayed 18
AC004 (V002, V001) (1,1,2) 154 133 on-time 13
AC015 (V002, V001) (1,1,2) 65 194 delayed 20
AC013 (V002, V006) (1,4,3) 55 147 on-time 16
AC018 (V002, V007) (1,2,3) 103 165 on-time 8
AC005 (V003, V002) (1,5,1) 83 177 on-time 4
AC006 (V005, V007) (4,3,3) 199 148 on-time 15
AC007 (V003, V002) (1,5,1) 100 183 on-time 5
AC008 (V007, V001) (2,3,2) 104 155 on-time 10
AC009 (V001, V005) (5,1,2) 67 189 delayed 17
AC010 (V001, V005) (5,1,2) 114 163 on-time 3
AC012 (V005, V001) (4,3,2) 90 124 on-time 12
AC014 (V001, V002) (5,2,1) 64 174 on-time 9
AC016 (V007, V005) (2,5,2) 109 189 on-time 14
AC017 (V004, V006) (5,3,3) 155 149 on-time 11
AC019 (V004, V002) (5,5,2) 104 147 on-time 1
AC020 (V003, V006) (1,2,3) 96 146 on-time 2
@@ 7 Conclusions
In this work, we introduce a novel mechanism that enables service providers to allocate on-demand requests from Advanced Air Mobility (AAM) vehicles, each with heterogeneous private valuations, to a capacity-constrained airspace. We evaluate the effectiveness of our approach using a physically accurate urban air delivery dataset and hypothesized electric air-taxi services in Northern California.
Our findings suggest several promising directions for future research. First, deriving formal guarantees regarding the neighborhood within which the fixed point operator (FP) is contractive, based on AAM market parameters, would be crucial for establishing theoretical convergence guarantees for our algorithmic approach, which has shown strong performance in simulations. Additionally, exploring formal incentive compatibility and fairness guarantees within the proposed allocation mechanism presents an exciting area for further study.
@@ Appendix B Proof of Theoretical Results
Appendix C Table of Notations

### v2

@@ 6.1 Qualitative Analysis
[…]
6:    $p_{e}\leftarrow p_{e}+\beta\quad\forall e\in\mathcal{C}$triangleright Increase Price for Overcapacitated Goods
7:until $\mathcal{C}=\emptyset$
8:Return $\hat{\mathbf{x}}_{u\in U}$
We compare these approaches using the Toulouse example at a 50% capacity level and also consider a slightly less constrained case (60% capacity). The set of goods remains the same across all approaches. Our evaluation includes the number of agents that are never allocated (dropped agents), the number of delayed agents, the average delay duration for delayed agents, the total number of times agents are rebased (including agents that are rebased-once and rebased-twice), the number of agents that are rebased at least once, and the average number of times agents are rebased. Lower values are preferable across all metrics, and the lowest value for each case is shown in green and bolded.
For the 60% capacity case, both our approach and the Budget-based comparison allocate all agents, with our approach additionally resulting in fewer delayed and rebased aircraft. In the more constrained 50% case, our approach significantly reduces the number of unallocated agents. While the Budget-based approach results in fewer delayed aircraft and a lower number of rebased agents, minimizing the number of unallocated agents is the more desirable outcome. The Profit-based approach performs worse in both cases across almost every metric, further highlighting the benefits of using artificial currency. We attribute this to the high cumulative costs incurred when agents must bid on multiple goods in this setting. These results show that, for comparable betas, our approach can better handle rising congestion due to the lower numbers of unallocated and rebased agents.
6.2 Sensitivity Analysis
(a) Iterations vs. Number of UAVs
(b) Percent of Unallocated UAVs vs. Number of UAVs
Figure 8: Variation of the number of iterations and percent of unallocated agents with respect to the number of agents as we vary capacity constraints on the resources.
(a) Iterations vs. Number of Auctions (I)
(b) Iterations vs. Inner Loop Update Parameter in Algorithm 1 (N)
(c) Iterations vs. Tolerance in Algorithm 1
Figure 9: Variation in number of iterations of Algorithm 1 with different algorithmic parameters
In this subsection, we conduct numerical studies to evaluate the impact of key design parameters on the performance of our algorithmic approach.
In Fig. 8(a), we present the variation in the number of iterations of Algorithm 1 with respect to different numbers of agents under various capacity constraints. This scenario can be understood as a setting where there is only one auction window in which all agents simultaneously request their desired goods on the time-extended graph. We observe that an increase in the number of agents leads to a higher number of contested goods, which requires more iterations of Algorithm 1 to compute a fractional competitive equilibrium. At 100% capacity, there are no contested resources, and the algorithm requires only 8 iterations regardless of the number of agents. As the capacity decreases, the number of iterations increases due to a higher number of contested goods. In Fig. 8(b), we analyze the variation in the percentage of unallocated agents with respect to different numbers of agents under various capacity constraints. We observe that when the resource capacity is 100%, all UAVs receive their desired paths. However, as the capacity decreases, a greater number of agents remain unallocated. Furthermore, for any given capacity level, the percentage of unallocated agents increases as more agents participate in the auction.
In Fig. 9, we study the variation in the number of iterations of Algorithm 1 with respect to different market parameters. First, in Fig. 9(a), we examine how the number of iterations of Algorithm 1 changes as we vary the number of auctions in our receding horizon approach. As the number of auctions increases, the number of UAVs participating in each auction decreases, resulting in fewer contested goods. Consequently, Algorithm 1 converges more quickly. Next, in Fig. 9(b), we evaluate the impact of the number of inner loop updates (parameter $N$ in Algorithm 1). Recall, that the goal of Algorithm 1 is to emulate fixed point iteration (FP). Towards this goal, the role of inner loop updates in Algorithm 1 is to estimate $\lambda^{\dagger}(\omega^{(k)})$ for every update of $\omega^{(k)}$ in the outer loop. The results show that if $N$ is lower, the number of iterations of Algorithm 1 is higher, as the inner loop cannot estimate $\lambda^{\dagger}(\omega^{(k)})$ accurately. Consequently, increasing $N$ decreases the number of iterations up to a point. However, if we continue to increase $N$ past this point then the additional steps in the inner loop do not help in the convergence of fixed point iteration (FP) and instead cause the iterations of Algorithm 1 to start increasing again. Finally, in Fig. 9(c), we analyze how the number of iterations varies with different convergence tolerances. For clear presentation, we introduced the variable $\alpha$, which is a multiplier for the nominal value of tolerances (i.e., $\textsf{tol}_{\textsf{CE}},\textsf{tol}_{\textsf{ICE}},\textsf{tol}_{\textsf{EAE}}$) described in our setup above and we study the variation in the number of iterations with respect of $\alpha.$ As expected, stricter tolerances (i.e., lower $\alpha$) generally increase the number of iterations, since the algorithm must satisfy more stringent stopping conditions.
7 Limitations
Here, we discuss some limitations of the current modeling framework. Addressing these limitations presents interesting directions for future research.
@@ 7 Limitations
    We assume that the AAM vehicles are not malicious and they act according to Algorithms 1 and 2. One way to implement our mechanisms is that before takeoff, each AAM vehicle interacts with the SP through a “proxy agent”, which is an autonomous entity that bids on its behalf based on Algorithm 1 and 2. The use of proxy agents is a popular method for implementing iterative auctions (Cramton et al., 2006b). Developing suitable auditing mechanisms to ensure that agents act honestly is an interesting open problem.
@@
    We consider that the communication between the service provider and each UAV is delay-free and noiseless. Once the vehicle has taken off, its route is fixed, and no further in-flight communication is required.
@@
    We assume that a single service provider controls the airspace. An interesting direction for future research is to extend our approach when the airspace is managed by multiple service providers.
@@
    In the current framework, each UAV is allocated to a feasible trajectory before departing and they do not change the trajectory during flight.
  • (v)
    We assume that adequate infrastructure and operational measures are in place to handle emergencies.
  • (vi)
    We consider that the SP randomly assigns budget to all AAM vehicles, but developing better budget allocation schemes that result in socially optimal outcome is an interesting direction of future research.
@@ 8 Conclusions
In this work, we introduce a novel mechanism that enables service providers to allocate on-demand requests from Advanced Air Mobility (AAM) vehicles, each with heterogeneous private valuations, to a capacity-constrained airspace. We evaluated the effectiveness of our approach using a physically accurate urban air delivery dataset. This is the first work in the AAM literature to allocate constrained airspace resources to dynamically arriving AAM vehicles without requiring knowledge of their private valuations. Core to our approach is an artificial currency-based auction mechanism that is implemented in a receding-horizon manner, wherein for each auction, we use a distributed iterative algorithm that accounts for individual agent preferences while ensuring system efficiency and safety.
@@ Appendix B Proof of Theoretical Results
Appendix C Derivation of Inner Loop Updates in Algorithm 1
The updates in the inner loop in Algorithm 1 is derived based on ADMM updates for (4). We review the basic structure of the ADMM algorithm in Section C.1 and then derive the inner loop updates in Section C.2.
C.1 Review of ADMM algorithm
The Alternative Direct Method of Multipliers (ADMM) is a distributed convex optimization algorithm that decomposes a problem into smaller subproblems, solves them in parallel, and coordinates to find a global solution via dual updates (Boyles et al., 2010; Boyd and Vandenberghe, 2004). It is built on dual ascent and augmented lagrangian methods.
Consider the following optimization problem with separable cost structure:
$\max_{\mathbf{x}\in X,\mathbf{y}\in Y}$ $\ h(\mathbf{x},\mathbf{y})=h_{1}(\mathbf{x})+h_{2}(\mathbf{y})$ (21)
s.t. $\ A\mathbf{x}+B\mathbf{y}=\mathbf{c},$
where
  • (i)
    $X\subset\mathbb{R}^{a},Y\subset\mathbb{R}^{b}$ are closed convex sets,
  • (ii)
    $h_{1}:\mathbb{R}^{a}\rightarrow\mathbb{R},h_{2}:\mathbb{R}^{b}\rightarrow\mathbb{R},$
  • (iii)
    $A\in\mathbb{R}^{s\times a},B\in\mathbb{R}^{s\times b},\mathbf{c}\in\mathbb{R}^{s}.$
Let $\mu\in\mathbb{R}^{s}$ be the dual multiplier of constraint in (21). Consider the following augmented Lagrangian function for (21) for some parameter $\beta>0$
$L_{\beta}(\mathbf{x},\mathbf{y})=h_{1}(\mathbf{x})+h_{2}(\mathbf{y})-\mu^{\top}(A\mathbf{x}+B\mathbf{y}-\mathbf{c})-\frac{\beta}{2}\|A\mathbf{x}+B\mathbf{y}-\mathbf{c}\|^{2}.$
The ADMM algorithm is a discrete-time algorithm, indexed by $k$, given as follows
$\mathbf{x}^{(n+1)}$ $=\arg\max_{\mathbf{x}\in X}L_{\beta}(\mathbf{x},\mathbf{y}^{(n)})$ (22)
$\mathbf{y}^{(n+1)}$ $=\arg\max_{\mathbf{y}\in Y}L_{\beta}(\mathbf{x}^{(n+1)},\mathbf{y})$
$\mu^{(n+1)}$ $=\mu^{(n)}+\beta(A\mathbf{x}^{(n+1)}+B\mathbf{y}^{(n+1)}-\mathbf{c}).$
The parameter $\beta$ is also referred to as the step-size parameter for the ADMM algorithm.
C.2 ADMM Updates for (4)
The inner loop in Algorithm 1 is nothing but the ADMM algorithm applied to (4).
For any $\beta>0$, we form the augmented Lagrangian $\mathcal{L}(\bar{\mathbf{x}},\mathbf{y},\mathbf{z},\lambda,\mathbf{p},\tilde{\mathbf{p}})$ for (4) as follows
$L_{\beta}(\bar{\mathbf{x}},\mathbf{y},\mathbf{z},\lambda,\mathbf{p},\tilde{\mathbf{p}})=$ $\sum_{u\in U}(w_{u}+\omega_{u})\log(f_{u}(\bar{\mathbf{x}}_{u}))-\sum_{u\in U}p_{\textsf{o}}x_{u,\textsf{o}}-\sum_{u\in U}\tilde{\mathbf{p}}_{u}^{\top}(\mathbf{x}_{u}-\mathbf{y}_{u})-\mathbf{p}^{\top}\left(\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\right)$
$-\sum_{u\in U}\lambda_{u}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)-\frac{\beta}{2}\sum_{u\in U}\|\mathbf{x}_{u}-\mathbf{y}_{u}\|^{2}-\frac{\beta}{2}\left\|\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\right\|^{2}.$ (23)
The ADMM algorithm (as per (22)) are given as follows:
[…]
$\hskip 85.35826pt-\sum_{u\in U}\lambda_{u}^{(n)}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)-\frac{\beta}{2}\sum_{u\in U}\|\mathbf{x}_{u}-\mathbf{y}_{u}^{(n)}\|^{2}$ (24a)
$({\mathbf{y}}^{(n+1)},\mathbf{z}^{(n+1)})$ $=\operatorname*{arg\,max}_{{\mathbf{y}}\in\mathbb{R}^{U\times|\tilde{\mathcal{E}}|},\ {\mathbf{z}}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}_{+}}L_{\beta}(\bar{\mathbf{x}}^{(n+1)},\mathbf{y},\mathbf{z},\lambda^{(n)},\mathbf{p}^{(n)},\tilde{\mathbf{p}}^{(n)})$
$=\operatorname*{arg\,max}_{{\mathbf{y}}\in\mathbb{R}^{U\times|\tilde{\mathcal{E}}|},\ {\mathbf{z}}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}_{+}}-\sum_{u\in U}\tilde{\mathbf{p}}_{u}^{(n)}{}^{\top}(\mathbf{x}_{u}^{(n+1)}-\mathbf{y}_{u})-\mathbf{p}^{(n)}{}^{\top}\left(\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\right)$
$\hskip 85.35826pt-\frac{\beta}{2}\sum_{u\in U}\|\mathbf{x}_{u}^{(n+1)}-\mathbf{y}_{u}\|^{2}-\frac{\beta}{2}\left\|\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\right\|^{2}$ (24b)
$\lambda^{(n+1)}_{u}$ $=\lambda_{u}+\beta(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}^{(n+1)}+x_{u,\varnothing}^{(n+1)}-1),\quad\forall\ u\in U$ (24c)
$\mathbf{p}^{(n+1)}$ $=\mathbf{p}^{(n)}_{u}+\beta(\sum_{u\in U}\mathbf{y}_{u}^{(n+1)}+\mathbf{z}^{(n+1)}-\ell)$ (24d)
$\tilde{\mathbf{p}}^{(n+1)}_{u}$ $=\tilde{\mathbf{p}}^{(n)}_{u}+\beta(\mathbf{x}_{u}^{(n+1)}-\mathbf{y}_{u}^{(n+1)}),\quad\forall\ u\in U.$ (24e)
First, we claim the if $\mathbf{p}^{(0)}=\tilde{\mathbf{p}}_{u}^{(0)}$ for every $u\in U$, then $\mathbf{p}^{(n)}=\tilde{p}_{u}^{(n)}$ for every $u\in U$ and $n\in\mathbb{N}$. We prove this by induction. Suppose for some $n,$ $\mathbf{p}^{(n)}=\tilde{p}_{u}^{(n)}$ for every $u\in U$ then we show that $\mathbf{p}^{(n+1)}=\tilde{p}_{u}^{(n+1)}$ for every $u\in U.$ To see this, note from the first order conditions of optimality for (24b) with respect to $\mathbf{y}$, we obtain
$\tilde{\mathbf{p}}_{u}^{(n)}-\mathbf{p}_{u}^{(n)}+\beta(\mathbf{x}_{u}^{(n+1)}-\mathbf{y}^{(n+1)})-\beta(\sum_{u\in U}\mathbf{y}_{u}^{(n+1)}+\mathbf{z}^{(n+1)}-\ell)=0.$ (25)
Using preceding equation, we obtain
[…]
This concludes our claim. Therefore, we get rid of notation $\tilde{\mathbf{p}}$ and work only with $\mathbf{p}.$
Finally, note that (24a) is separable in $\bar{\mathbf{x}}_{u}$ for every $u\in U$. Against the preceding backdrop, (24) can be re-written as
[…]
$\hskip 85.35826pt-\lambda_{u}^{(n)}(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}+x_{u,\varnothing}-1)-\frac{\beta}{2}\|\mathbf{x}_{u}-\mathbf{y}_{u}^{(n)}\|^{2}$ (26a)
$({\mathbf{y}}^{(n+1)},\mathbf{z}^{(n+1)})$ $=\operatorname*{arg\,max}_{{\mathbf{y}}\in\mathbb{R}^{U\times|\tilde{\mathcal{E}}|},\ {\mathbf{z}}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}_{+}}-\mathbf{p}^{(n)}{}^{\top}\mathbf{z}-\frac{\beta}{2}\sum_{u\in U}\|\mathbf{x}_{u}^{(n+1)}-\mathbf{y}_{u}\|^{2}-\frac{\beta}{2}\left\|\sum_{u\in U}\mathbf{y}_{u}+\mathbf{z}-\ell\right\|^{2}$ (26b)
$\lambda^{(n+1)}_{u}$ $=\lambda_{u}+\beta(\tilde{\mathbf{a}}_{u}^{\top}\mathbf{x}_{u}^{(n+1)}+x_{u,\varnothing}^{(n+1)}-1),\quad\forall\ u\in U$ (26c)
$\mathbf{p}^{(n+1)}$ $=\mathbf{p}^{(n)}_{u}+\beta(\sum_{u\in U}\mathbf{y}_{u}^{(n+1)}+\mathbf{z}^{(n+1)}-\ell)$ (26d)
Updates (26) correspond to the inner loop updates in Algorithm 1, where (26a) is implemented locally by different AAM vehicles and (26b)-(26d) are implemented by service provider.
Appendix D Vertiport Reservation Mechanism in Northern California
Figure 11: Northern California Vertiport Map. This map, adapted from a Google Maps image, highlights seven distinct vertiports using unique color codes and displays the example routes as red lines.
In this example, 20 air taxis request a departure, air route, and landing clearances among seven vertiport destinations during a 10-minute auction window. The requests by the air taxis, the final allocation of routes, and payments to the SP are presented in Table 2 along with the maximum capacity in every segment of the desired routes, the utility of the air taxis for a given path, and their initial air credits. We set $\beta=50,p_{\textsf{o}}=10,v_{u,\textsf{o}}=1,v_{u,\varnothing}=1,N=2,\textsf{tol}=1\times 10^{-4}$. The Maximum Capacity column in Table 2 specifies the maximum number of vehicles that can traverse a travel segment at any given time. These values help the reader identify contested travel segments and understand why certain agents must compete for access. In these tables, air taxis sharing the same color represent those that simultaneously requested the same trajectory slot, leading to a constraint violation. As a result, only a subset of these air taxis were granted their preferred route, while the others were denied access. We also present the rank number of these agents representing the order in which each agent computed their integral allocation, as outlined in Algorithm 2.
Below, we highlight the main observations from our numerical study.
  • (i)
  • (ii)
  • (iii)
  • (iv)
Table 2: Results of the allocation of air taxis to the desired routes, payments to the SP, utility, initial air credits, and maximum capacity in the en-route travel segment
Aircraft Req. Route (Orig., Dest.) Req. Time (Arr, Dep) Max. Capacity (Dep, Route, Arr) Allocated Time (Arr, Dep) Status Price (\$) Initial Air Credits Utility Rank
AC001 (V007, V002) (16, 54) (2,4,1) (16, 54) on-time 0.0 125 118 6
AC002 (V005, V004) (19, 47) (4,5,1) (19, 47) on-time 5.73 90 171 7
AC011 (V006, V004) (19, 47) (1,2,1) (20, 48) delayed 1.53 78 135 19
AC003 (V002, V001) (16, 21) (1,1,2) (19, 24) delayed 3.71 135 172 18
AC004 (V002, V001) (16, 21) (1,1,2) (16, 21) on-time 20.36 154 133 13
AC015 (V002, V001) (16, 21) (1,1,2) (20, 25) delayed 0.86 65 194 20
AC013 (V002, V006) (17, 41) (1,4,3) (17, 41) on-time 11.11 55 147 16
AC018 (V002, V007) (18, 56) (1,2,3) (18, 56) on-time 8.46 103 165 8
AC005 (V003, V002) (11, 19) (1,5,1) (11, 19) on-time 0.0 83 177 4
AC006 (V005, V007) (18, 68) (4,3,3) (18, 68) on-time 0.0 199 148 15
AC007 (V003, V002) (15, 23) (1,5,1) (15, 23) on-time 0.0 100 183 5
AC008 (V007, V001) (12, 54) (2,3,2) (12, 54) on-time 0.0 104 155 10
AC009 (V001, V005) (13, 34) (5,1,2) (14, 35) delayed 2.10 67 189 17
AC010 (V001, V005) (13, 34) (5,1,2) (13, 34) on-time 5.75 114 163 3
AC012 (V005, V001) (16, 37) (4,3,2) (16, 37) on-time 0.0 90 124 12
AC014 (V001, V002) (11, 24) (5,2,1) (11, 24) on-time 0.0 64 174 9
AC016 (V007, V005) (17, 67) (2,5,2) (17, 67) on-time 0.0 109 189 14
AC017 (V004, V006) (16, 44) (5,3,3) (16, 44) on-time 0.0 155 149 11
AC019 (V004, V002) (16, 35) (5,5,2) (16, 35) on-time 0.0 104 147 1
AC020 (V003, V006) (16, 38) (1,2,3) (16, 38) on-time 0.11 96 146 2
Appendix E Table of Notations